OpenStudy (anonymous):

Solve the Proportion. 3/5 = 6/x + 3 @dc104

OpenStudy (anonymous):

This question brought a variable in, so automatically, I don't know if the method will work.

OpenStudy (anonymous):

Start by multiplying both sides by x.

OpenStudy (anonymous):

3/5 - 3 = 6/x (3/5 - 3)*x = 6 x = 6/(3/5 - 3) x = -5/2

OpenStudy (anonymous):

In this case, the only difference is I have moved the 3 over to the other side first. Then I proceed with the process I explained to you before

OpenStudy (anonymous):

Okay. I have one more question I need answering. Otherwise I'll get through the remainder. If I need help, I know who to call. @dc104

OpenStudy (anonymous):

Do you want to just answer it here?

OpenStudy (anonymous):

No problem